OpenStudy (anonymous):

Which system has (-8, 0) as the solution Which system has infinite solutions? Which system has (-2, -2) as the solution Which system has (1, 3) as the solution Which system has no solutions? A) 4x + 6y = -12 2x + 3y = 9 B) x + y = -4 -3x + 2y = 2 C) x + y = -1 4x + 4y = -4 D) x - 4y = -8 2x - 3y = -16 E) x - y = -2 3x - y = 0 Can someone please help? I already solved all these but every time I graph them I cant find where they intersect

OpenStudy (solomonzelman):

4x + 6y = -12 2x + 3y = 9 ----> 4x+6y=18 subtract them from each other 0= -30 A is no solution.

OpenStudy (solomonzelman):

x + y = -4 ----> 3x+3y=-12 -3x + 2y = 2 add them to each other, 5y=-10 y=-2. I can see that B has (-2, -2) as the solution

OpenStudy (solomonzelman):

x + y = -1 ---> 4x+4y=-4 4x + 4y = -4 C has infinite solutions

OpenStudy (anonymous):

I thought I had to graph them

OpenStudy (solomonzelman):

Hold on, first lets see which one is which.

OpenStudy (anonymous):

Okay, thank you for helping by the way

OpenStudy (solomonzelman):

x - 4y = -8 ----> 2x-8y= -16 2x - 3y = -16 subtract them from each other, -5y=0 D has (-8, 0) as the solution

OpenStudy (solomonzelman):

Guess which one is E.

OpenStudy (anonymous):

(1,3) (:

OpenStudy (solomonzelman):

A no solution B has (-2, -2) as the solution C has infinite solutions D has (-8, 0) as the solution E has (1,3) as the solution to get this straight.

OpenStudy (solomonzelman):

You don't have to graph them, the question doesn't tell you too, but if you want I can help you graph.

OpenStudy (anonymous):

Yes, thank you(: they were all right but I was just trying to graph them so that I could figure out where they intersect but I understand graphing already, thanks for offering though!

OpenStudy (solomonzelman):

Anytime, I wouldn't what to graph them either, so it is not a problem that you refused too. Less work for both of us.

OpenStudy (anonymous):

Yeah I didn't want to make you do that, its a lot of work and a waste of our time since I already understand it lol

OpenStudy (solomonzelman):

Cool! \[(•_‿•) \]